You have been asked to develop a capitation rate for a primary care group based on the following projections: Service Annual Frequency/1,000 cost per service inpatient visits 100 $7000 office visits 3,000 $45.00 lab/x-ray 500 $25.00

What per-member per-month (PMPM) rate would be required to break even, ignoring any copayment?
Answer #1

Solution

Step 1:

Calculate per patient Utilization rate:

Inpatient utilization rate = Total inpatient visit / total members

= 100 / 1000

= 0.1 inpatient visit per member in a year

Office utilization rate = Total office utilization rate / total members

= 3000 / 1000

= 3.0 visit per member in a year.

Lab utilization rate = Total lab utilization rate / Total member

= 500 / 1000

= 0.5 visit per member in a year.

Step 2: Calculate per-member per-month

PMPM = Utilization rate X cost / 12

PMPM for inpatient visits = 0.1 X $ 7000 / 12

= $58.33

PMPM for office visits = 3.0 X $45 / 12

= $11.25

PMPM for lab visits = 0.5 X $25 / 12

= $1.04

Step 3:

Per-member per-month rate would be $58.33, $11.25, $1.04 for the inpatient visits, office visits, and lab/x-ray visits respectively.